Use the calculus to determine the approximate change in the area of a circle when the radius increases from 3 cm to 3.02 cm.
Answer by Alan3354(69443)

Use the calculus to determine the approximate change in the area of a circle when the radius increases from 3 cm to 3.02 cm.
=================
Area = pi*r^2
dArea = 2pi*r*dr
dArea = 2pi*3*0.02 = 0.36799
------------------
pi*3.02^2 - pi*9 = 0.37825